What is the probability that 9 package recipients are unemployed


1
Expert's answer
2021-05-07T13:31:02-0400

Probability of unemployed people p = 0.11


Then "q=1-p=1-0.11=0.89"

"n=10"


Probability that 9 package recipients are unemployed "= ^{10}C_9p^9q^1"


"=10\times (0.11)^9(0.89)^1=0.00002098"